# Offline Model-Based Optimization: Comprehensive Review

### v1

@@ 5.3 Autoregressive Model
Autoregressive models are widely adopted for generative tasks, particularly in language modeling. Notable examples include LSTMs (Hochreiter, 1997) and Transformer-based models (Vaswani et al., 2017; Brown et al., 2020), which factorize the joint distribution in an autoregressive manner. Specifically, the joint probability of a sequence is expressed as:
$p_{\bm{\theta}}(\bm{x})=\prod_{i=1}^{N}p_{\bm{\theta}}\left(\bm{x}_{i}\mid\bm{x}_{<i}\right),$
and the model is typically trained by maximizing the log-likelihood, often using the cross-entropy loss for discrete tokens. During generation, elements are sampled sequentially, with each new token conditioned on the previously generated tokens.
@@ 5.4 Diffusion Model
Diffusion models, a subset of latent variable models, gradually perturb data by injecting Gaussian noise during the forward process. The reverse process iteratively denoises the data using a learned score estimator (Ho et al., 2020). These models can be represented in a continuous framework via a stochastic differential equation (SDE), as discussed in Song et al. (2021). The forward SDE is expressed as:
$d\bm{x}=\bm{h}(\bm{x},t)\,dt+g(t)\,d\bm{w},$ (10)
where $\bm{h}(\cdot,t):\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ denotes the drift coefficient, $g(t):\mathbb{R}\rightarrow\mathbb{R}$ is the diffusion coefficient, and $\bm{w}$ represents a standard Wiener process. This formulation progressively converts the data distribution into a noise distribution. The reverse process is characterized by:
$d\bm{x}=\left[\bm{h}(\bm{x},t)-g(t)^{2}\nabla_{\bm{x}}\log p_{t}(\bm{x})\right]\,dt+g(t)\,d\bar{\bm{w}},$ (11)
with $\nabla_{\bm{x}}\log p_{t}(\bm{x})$ indicating the score of the marginal distribution at time $t$, and $\bar{\bm{w}}$ symbolizing the reverse-time Wiener process. The score function is approximated using a time-dependent neural network $\bm{s}_{\bm{\theta}}(\bm{x}_{t},t)$, which facilitates the transformation of noise back into samples. They have demonstrated exceptional performance in synthesizing high-fidelity images and a wide range of complex data types.
@@
In inverse methods, the conditional diffusion model is trained using either vanilla guidance, where the label or condition is directly provided as an input (Chang et al., 2023), or classifier-free guidance, which derives guidance by contrasting the outputs of a conditional model with those of an unconditional model (Ho & Salimans, 2022). For example, Zhang et al. (2024) proposes a vanilla guidance approach that learns a weight function to assign higher weights to high-performing designs. This method focuses the training of the diffusion model $\bm{s}_{\theta}(\bm{x}_{t},y_{c})$ on high-performing designs, using them as implicit high-performing conditions $y_{c}$ during sampling. However, vanilla guidance lacks an adjustable parameter to control sampling strength, which motivates the use of classifier-free guidance where the strength can be tuned via a parameter $\omega$. The corresponding score function is defined as follows:
@@
In this context, Lee et al. (2023) investigate guided molecule generation toward high-performing regions with respect to target properties such as protein-ligand interactions, drug-likeness, and synthesizability. In Yuan et al. (2024b), gradient ascent is first employed to optimize the design. To address potential out-of-distribution issues, the method subsequently recovers the corresponding latent representation by injecting diffusion noise and then applying a denoising procedure, yielding a sample that conforms to the diffusion prior. This approach can be interpreted as a variant of classifier guidance due to its use of classifier gradient. Compared to classifier-free guidance, classifier guidance is less frequently adopted, likely due to the additional training cost of an extra surrogate model and the potential risk of reward hacking associated with the classifier.
@@ 5.5 Flow Matching
Flow matching defines a time-dependent conditional probability trajectory $p_{t}(\bm{x}\mid\bm{x}_{1})$ for $t\in[0,1]$, starting from $p_{0}(\bm{x}\mid\bm{x}_{1})=q(\bm{x})$ and evolving toward an approximate Dirac delta, $p_{1}(\bm{x}\mid\bm{x}_{1})\approx\delta(\bm{x}-\bm{x}_{1})$, where $\bm{x}_{1}$ is drawn from $p_{\text{data}}$. This evolution is conditioned on a specific data sample $\bm{x}_{1}$ and driven by a time-dependent vector field $u_{t}(\bm{x}\mid\bm{x}_{1})$. A neural network with parameters $\bm{\theta}$ is used to learn the marginal vector field,
$\hat{v}(\bm{x},t;\theta)\approx v(\bm{x},t)=\mathbb{E}_{\bm{x}_{1}\sim p_{t}(\bm{x}_{1}\mid\bm{x})}\Bigl{[}u_{t}(\bm{x}\mid\bm{x}_{1})\Bigr{]},$ (14)
which acts as a neural ODE that transforms the noise distribution $q(\bm{x})$ into the data distribution $p_{\text{data}}(\bm{x})$.
Following Pooladian et al. (2023), the process begins by sampling a noise vector $\bm{x}_{0}$ from $q(\bm{x}_{0})$ and linearly interpolating it with a data point $\bm{x}_{1}$ via $\bm{x}\mid\bm{x}_{1},t=(1-t)\bm{x}_{0}+t\bm{x}_{1}$ (with $\bm{x}_{0}\sim q(\bm{x}_{0})$). The corresponding conditional vector field is defined as $u_{t}(\bm{x}\mid\bm{x}_{1})=\frac{\bm{x}_{1}-\bm{x}}{1-t}$, which, given the interpolation, simplifies to $u_{t}(\bm{x}\mid\bm{x}_{1})=\bm{x}_{1}-\bm{x}_{0}$. The training objective minimizes the loss
$\mathcal{L}(\theta)=\mathbb{E}_{t,\,p_{\text{data}}(\bm{x}_{1}),\,q(\bm{x}_{0})}\left\|\hat{v}(\bm{x},t;\theta)-(\bm{x}_{1}-\bm{x}_{0})\right\|^{2}.$ (15)
Once trained, the vector field $\hat{v}(\bm{x},t;\theta)$ is used to generate new samples by integrating the neural ODE.
@@
Because flow matching closely resembles diffusion models, analogous conditional generation techniques can be applied. In particular, both inverse method classifier-free guidance (Zheng et al., 2023) and forward method classifier guidance (Dao et al., 2023) are readily adaptable within the flow matching framework. Regarding the inverse method, Stärk et al. (2024) introduce Dirichlet flow matching on the simplex and extend classifier-free guidance to more effectively steer the sequence generation process. In the forward method, Yuan et al. (2024a) investigate the use of multiple surrogate models to guide flow sampling toward the Pareto-front in multi-objective optimization settings. Moreover, Chen et al. (2025) propose training an affinity predictor to steer protein conformation sampling toward stable configurations within the AlphaFlow framework (Jing et al., 2024).
@@ 5.6 Energy-Based Model
where $Z(\theta)$ denotes the partition function. Training EBMs typically aims to assign lower energy to observed data while raising the energy of samples drawn from the model. A common objective is contrastive divergence:
$\mathcal{L}_{\mathrm{EBM}}=\mathbb{E}_{\bm{x}\sim p_{\text{data}}(\bm{x})}\left[E_{\theta}(\bm{x})\right]-\mathbb{E}_{\bm{x}\sim p_{\theta}(\bm{x})}\left[E_{\theta}(\bm{x})\right].$
This formulation encourages the model to distinguish between real data and generated (negative) samples.
Sampling from an EBM is challenging due to the intractability of the partition function, and it generally relies on Markov Chain Monte Carlo (MCMC) techniques (Neal, 1993; Hinton, 2002; Tieleman, 2008). A widely used approach is Langevin dynamics (Welling & Teh, 2011), with the iterative update:
$\bm{x}_{k+1}=\bm{x}_{k}-\frac{\epsilon}{2}\nabla_{\bm{x}}E_{\theta}(\bm{x}_{k})+\sqrt{\epsilon}\,\bm{\eta}_{k},\quad\bm{\eta}_{k}\sim\mathcal{N}(0,I),$
where $\epsilon$ is the step size. This stochastic process, which introduces Gaussian noise $\bm{\eta}_{k}$, is designed to asymptotically sample from $p_{\theta}(\bm{x})$. Often, hundreds to thousands of iterations are needed to ensure adequate mixing, and slow mixing remains a notable challenge, especially in high-dimensional spaces.
EBMs provide a flexible framework for modeling complex data distributions by directly parameterizing an energy function. While they share a fundamental relation with score-based models (Song et al., 2021)—since $\nabla_{\bm{x}}\log p_{\theta}(\bm{x})=-\nabla_{\bm{x}}E_{\theta}(\bm{x})$—the two approaches diverge in practice: EBMs require explicit energy modeling and rely on MCMC-based sampling (e.g., via Langevin dynamics), whereas score-based models aim to directly learn the score function without necessarily specifying an energy function. We note that GFlowNets (Bengio et al., 2023) can amortize the MCMC process, making them applicable to EBMs (Zhang et al., 2022). Although these methods enable faster mode mixing to estimate the partition function to compare with MCMC methods, thus improving the practicality of EBMs, none of these techniques have been applied to offline optimization. We believe that exploring this direction could offer promising avenues for future research.
@@
It is important to note that the distinction between forward and inverse methods in EBMs is often subtle, as both the energy function and a surrogate model essentially map a design to a scalar value. In Beckham & Pal (2023), the authors reinterpret the original forward method, COMs, as an EBM trained via contrastive divergence. COMs optimizes two losses—a mean-squared error loss for surrogate modeling and a conservative objective for the EBM—using a shared network architecture, where the resulting energy function steers the sampling process towards high-performing designs, a characteristic typically associated with inverse methods. Moreover, Beckham & Pal (2023) further proposes a decoupled version of COMs, in which separate networks are employed for the surrogate and the EBM, reinforcing its classification as a forward method.
@@ 5.7 Control by Generative Flow Network (GFlowNet)
Unlike earlier sections that focus on general generative models and their conditional generation approaches, this section introduces a control method for these models: Generative Flow Networks  (GFlowNets), which have recently gained popularity. GFlowNets (Bengio et al., 2021; 2023) model the generative process as a sequential decision-making problem, where a complete candidate solution $\bm{x}$ is constructed through a sequence of transitions. Their goal is to perform amortized inference by sampling from a distribution proportional to a reward function, i.e., $p(\bm{x})\propto R(\bm{x})$. This approach produces samples that are both high-reward and diverse, which is particularly beneficial when the reward function is an imperfect surrogate learned from an offline dataset. In contrast, standard RL methods that simply maximize $R(\bm{x})$ can be unsafe under epistemic uncertainty, making GFlowNets especially attractive in offline optimization scenarios.
It is important to note that GFlowNets represent a learning method rather than a standalone generative model. Their sequential decision-making process is typically implemented by generative models such as autoregressive models, diffusion models, or any model that provides sequential inference structures.
GFlowNets Contraints
In GFlowNets, a sample $\bm{x}$ is generated by traversing a directed acyclic graph from an initial state $\bm{s}_{0}$ to a terminal state $\bm{s}_{T}=\bm{x}$ through a sequence of transitions:
$\bm{s}_{0}\;\to\;\bm{s}_{1}\;\to\;\cdots\;\to\;\bm{s}_{T}=\bm{x}.$
The central training criterion enforces flow consistency: for each intermediate state $\bm{s}^{\prime}$, the total incoming flow must equal the total outgoing flow:
$\sum_{\bm{s}\in\mathrm{Parents}(\bm{s}^{\prime})}F(\bm{s};\theta)\,P_{F}\!\bigl{(}\bm{s}^{\prime}\mid\bm{s};\theta\bigr{)}\;=\;F(\bm{s}^{\prime};\theta).$
At the terminal state, the flow is fixed to the reward,
$F(\bm{s}_{T};\theta)\;=\;R(\bm{x}),$
which serves as the surrogate objective in this work.
A common way to implement flow consistency is via the detailed balance (DB) condition, introducing a backward policy $P_{B}(\bm{s}\mid\bm{s}^{\prime};\theta)$ to match flows:
$F(\bm{s};\theta)\,P_{F}\!\bigl{(}\bm{s}^{\prime}\mid\bm{s};\theta\bigr{)}\;=\;F(\bm{s}^{\prime};\theta)\,P_{B}\!\bigl{(}\bm{s}\mid\bm{s}^{\prime};\theta\bigr{)},\quad\bm{s}^{\prime}\in\mathrm{Children}(\bm{s}).$
This one-step condition is similar to temporal-difference (TD) learning in RL. It can be extended to longer trajectories: a partial subsequence $\bm{s}_{m}\to\cdots\to\bm{s}_{l}$ (sub-trajectory balance, Sub-TB (Madan et al., 2023)) or a complete sequence $\bm{s}_{0}\to\cdots\to\bm{s}_{n}=\bm{x}$ (trajectory balance, TB (Malkin et al., 2022)). These extensions are analogous to TD-$\lambda$ and Monte Carlo learning methods. Sub-TB is formulated as:
$F(\bm{s}_{m};\theta)\,\prod_{i=m+1}^{l}P_{F}\!\bigl{(}\bm{s}_{i}\mid\bm{s}_{i-1};\theta\bigr{)}\;=\;F(\bm{s}_{l};\theta)\,\prod_{i=m+1}^{l}P_{B}\!\bigl{(}\bm{s}_{i-1}\mid\bm{s}_{i};\theta\bigr{)},$
while the TB condition is:
$Z_{\theta}\prod_{i=1}^{n}P_{F}\!\bigl{(}\bm{s}_{i}\mid\bm{s}_{i-1};\theta\bigr{)}\;=\;R(\bm{x})\prod_{i=1}^{n}P_{B}\!\bigl{(}\bm{s}_{i-1}\mid\bm{s}_{i};\theta\bigr{)},$
where $Z_{\theta}=F(\bm{s}_{0})$ is the partition function and $R(\bm{x})=F(\bm{s}_{n})$ is the terminal reward.
Recent work refines TB for posterior inference by introducing relative trajectory balance (RTB) (Venkatraman et al., 2024), where
$p_{\text{post}}(\bm{x})\;\propto\;p_{\text{prior}}(\bm{x})\,R(\bm{x}).$
In RTB, the backward policy $P_{B}$ is replaced by a prior distribution $P_{\text{prior}}$, yielding:
$Z_{\theta}\prod_{i=1}^{n}P_{\text{post}}\!\bigl{(}\bm{s}_{i}\mid\bm{s}_{i-1};\theta\bigr{)}\;=\;R(\bm{x})\prod_{i=1}^{n}P_{\text{prior}}\!\bigl{(}\bm{s}_{i}\mid\bm{s}_{i-1}\bigr{)}.$
This can be applied to active learning and offline optimization, allowing for a conservative strategy by restricting samples to the prior distribution learned from offline data, thus reducing risk under uncertainty.
Controlling Generative Models via GFlowNets
@@
For example, Jain et al. (2022) generates desirable biological sequences from scratch by defining a target reward function as a UCB score from a surrogate model. To improve robustness against the imperfections of the surrogate model at early stages, Kim et al. (2024a) conservatively search promising regions by introducing a parameter $\delta$, which is adjusted based on the prediction model’s uncertainty. Additionally, Ghari et al. (2023) utilize GFlowNets to modify existing sequences to enhance target properties, while Jain et al. (2023) employ conditional GFlowNets to generate diverse Pareto-optimal solutions for multi-objective optimization problems. Note that they use autoregressive models to produce these biological sequences and employ GFlowNets to control them effectively.
@@
One of the main challenges in offline optimization is the exploration-exploitation trade-off. Unlike RL, GFlowNets can effectively balance exploration and exploitation using a temperature parameter $\beta$, i.e., $p(\bm{x}|\beta)\propto R(\bm{x})^{\beta}$. Kim et al. (2024b) propose an effective way to learn such a policy by introducing a logit scaling network and verify that it achieves high extrapolation ability in offline optimization tasks. They use directional token generative models, which can generate sequences by prepending or appending tokens, and employ GFlowNets to control them.
@@
We can also formulate the offline optimization problem as posterior inference, where we have a pre-trained policy $p(\bm{x})$ and a reward function $R(\bm{x})$ trained on a given dataset. In this setting, we can use the RTB loss to obtain an unbiased sampler that amortizes the posterior distribution. For example, Yun et al. (2025) propose a posterior inference method with GFlowNets for diffusion models in black-box online optimization. This work may provide a key insight for applying posterior inference with GFlowNets to offline optimization.

### v2

@@ 5.3 Autoregressive Model
Autoregressive models are widely adopted for generative tasks, particularly in language modeling. Notable examples include LSTMs (Hochreiter, 1997) and Transformer-based models (Vaswani et al., 2017; Brown et al., 2020), which factorize the joint distribution in an autoregressive manner.
@@ 5.4 Diffusion Model
Diffusion models, a subset of latent variable models, gradually perturb data by injecting Gaussian noise during the forward process. The reverse process iteratively denoises the data using a learned score estimator (Ho et al., 2020). The score function is approximated using a time-dependent neural network $\bm{s}_{\bm{\theta}}(\bm{x}_{t},t)$, which facilitates the transformation of noise back into samples. They have demonstrated exceptional performance in synthesizing high-fidelity images and a wide range of complex data types.
@@
In inverse methods, the conditional diffusion model is trained using either vanilla guidance, where the label or condition is directly provided as an input (Chang et al., 2025), or classifier-free guidance, which derives guidance by contrasting the outputs of a conditional model with those of an unconditional model (Ho & Salimans, 2022). For example, Zhang et al. (2024) proposes a vanilla guidance approach that learns a weight function to assign higher weights to high-performing designs. This method focuses the training of the diffusion model $\bm{s}_{\theta}(\bm{x}_{t},y_{c})$ on high-performing designs, using them as implicit high-performing conditions $y_{c}$ during sampling. However, vanilla guidance lacks an adjustable parameter to control sampling strength, which motivates the use of classifier-free guidance where the strength can be tuned via a parameter $\omega$. The corresponding score function is defined as follows:
@@
In this context, Lee et al. (2023) investigate guided molecule generation toward high-performing regions with respect to target properties such as protein-ligand interactions, drug-likeness, and synthesizability. In Yuan et al. (2024), gradient ascent is first employed to optimize the design. To address potential out-of-distribution issues, the method subsequently recovers the corresponding latent representation by injecting diffusion noise and then applying a denoising procedure, yielding a sample that conforms to the diffusion prior. This approach can be interpreted as a variant of classifier guidance due to its use of classifier gradient. Compared to classifier-free guidance, classifier guidance is less frequently adopted, likely due to the additional training cost of an extra surrogate model and the potential risk of reward hacking associated with the classifier.
@@ 5.5 Flow Matching
Because flow matching closely resembles diffusion models, analogous conditional generation techniques can be applied. In particular, both inverse method classifier-free guidance (Zheng et al., 2023) and forward method classifier guidance (Dao et al., 2023) are readily adaptable within the flow matching framework. Regarding the inverse method, Stärk et al. (2024) introduce Dirichlet flow matching on the simplex and extend classifier-free guidance to more effectively steer the sequence generation process. In the forward method, Yuan et al. (2025) investigate the use of multiple surrogate models to guide flow sampling toward the Pareto-front in multi-objective optimization settings, with empirical results indicating superior performance over diffusion models (Yao et al., 2024b). Moreover, Chen et al. (2025) propose training an affinity predictor to steer protein conformation sampling toward stable configurations within the AlphaFlow framework (Jing et al., 2024).
@@ 5.6 Energy-Based Model
where $Z(\theta)$ denotes the partition function. Training EBMs typically aims to assign lower energy to observed data while raising the energy of samples drawn from the model. A common objective is contrastive divergence which encourages the model to distinguish between real data and generated (negative) samples. Sampling from an EBM is challenging due to the intractability of the partition function, and it generally relies on Markov Chain Monte Carlo (MCMC) techniques (Neal, 1993; Hinton, 2002; Tieleman, 2008).
We note that GFlowNets (Bengio et al., 2023) can amortize the MCMC process, making them applicable to EBMs (Zhang et al., 2022). Although these methods enable faster mode mixing to estimate the partition function to compare with MCMC methods, thus improving the practicality of EBMs, none of these techniques have been applied to offline black-box optimization. We believe that exploring this direction could offer promising avenues for future research.
@@
It is important to note that the distinction between forward and inverse methods in EBMs is often subtle, as both the energy function and a surrogate model essentially map a design to a scalar value. In Beckham & Pal (2023), the authors reinterpret the original forward method, conservative objective models (COMs), as an EBM trained via contrastive divergence. COMs optimizes two losses—a mean-squared error loss for surrogate modeling and a conservative objective for the EBM—using a shared network architecture, where the resulting energy function steers the sampling process towards high-performing designs, a characteristic typically associated with inverse methods. Moreover, Beckham & Pal (2023) further proposes a decoupled version of COMs, in which separate networks are employed for the surrogate and the EBM, reinforcing its classification as a forward method.
@@ 5.7 Control by Generative Flow Network (GFlowNet)
Unlike earlier sections on general generative models and their conditional approaches, this section introduces a control method: Generative Flow Networks (GFlowNets) (Bengio et al., 2021; 2023). GFlowNets model generation as a sequential decision process, where a solution $\bm{x}$ is constructed through a trajectory of transitions. Their goal is amortized inference—sampling from a distribution proportional to a reward function, $p(\bm{x})\propto R(\bm{x})$. This produces samples that are both high-reward and diverse, which is especially important when the reward is an imperfect surrogate learned from offline data. Unlike standard RL, which focuses on maximizing $R(\bm{x})$ and may be unsafe under epistemic uncertainty, GFlowNets balance exploration and exploitation, making them attractive in offline black-box optimization. A sample is generated by traversing a directed acyclic graph from an initial state $\bm{s}_{0}$ to a terminal state $\bm{s}_{T}=\bm{x}$:
$\bm{s}_{0}\to\bm{s}_{1}\to\cdots\to\bm{s}_{T}.$
Training is based on the flow consistency condition: for any intermediate state $\bm{s}^{\prime}$, incoming flow equals outgoing flow,
$\sum_{\bm{s}\in\mathrm{Parents}(\bm{s}^{\prime})}F(\bm{s};\theta)\,P_{F}(\bm{s}^{\prime}\mid\bm{s};\theta)=F(\bm{s}^{\prime};\theta),$
with terminal flow fixed to the reward,
$F(\bm{s}_{T};\theta)=R(\bm{x}).$
Different training objectives instantiate this principle, including trajectory balance (TB) (Malkin et al., 2022) and sub-trajectory balance (SubTB) (Madan et al., 2023). These objectives have been used to control diverse generative models, including molecular graphs (Bengio et al., 2021), bidirectional string models (Kim et al., 2024c), and diffusion models (Lahlou et al., 2023a).
Offline MBO applications
@@
For example, Jain et al. (2022) generates desirable biological sequences from scratch by defining a target reward function as the upper confidence bound score from a surrogate model. To improve robustness against the imperfections of the surrogate model at early stages, Kim et al. (2024a) conservatively search promising regions by introducing a parameter $\delta$, which is adjusted based on the prediction model’s uncertainty. Additionally, Ghari et al. (2023) utilize GFlowNets to modify existing sequences to enhance target properties, while Jain et al. (2023) employ conditional GFlowNets to generate diverse Pareto-optimal solutions for multi-objective optimization problems. Note that they use autoregressive models to produce these biological sequences and employ GFlowNets to control them effectively.
@@
One of the main challenges in offline black-box optimization is the exploration-exploitation trade-off. Unlike RL, GFlowNets can effectively balance exploration and exploitation using a temperature parameter $\beta$, i.e., $p(\bm{x}|\beta)\propto R(\bm{x})^{\beta}$. Kim et al. (2024b) propose an effective way to learn such a policy by introducing a logit scaling network and verify that it achieves high extrapolation ability in offline black-box optimization tasks. They use directional token generative models, which can generate sequences by prepending or appending tokens, and employ GFlowNets to control them.
@@
We can also formulate the offline black-box optimization problem as posterior inference, where we have a pre-trained policy $p(\bm{x})$ and a reward function $R(\bm{x})$ trained on a given dataset. In this setting, we can use the RTB loss to obtain an unbiased sampler that amortizes the posterior distribution. For example, Yun et al. (2025) propose a posterior inference method with GFlowNets for diffusion models in black-box online black-box optimization. This work may provide a key insight for applying posterior inference with GFlowNets to offline black-box optimization.
Comparison of Generative Models in Offline MBO.
We have presented a range of generative models used in offline MBO. Below, we summarize their respective strengths and limitations in this context.
  • Variational autoencoders provide a continuous, differentiable latent space suitable for gradient-based offline MBO, but they can suffer from posterior collapse and reconstruction mismatch that may lead to inaccurate property estimates.
  • Generative adversarial networks generate high-quality samples but typically lack an explicit latent space for controllable optimization, often requiring inverse methods. Additionally, training can be unstable due to mode collapse. The GAN discriminator is frequently used to detect out-of-distribution (OOD) designs.
  • Autoregressive models are well-suited for sequence generation and are often combined with reinforcement learning techniques for guided sampling, though the sequential sampling process is inherently slow.
  • Diffusion models provide stable training, high sample fidelity, and natural diversity. They are computationally intensive at inference time due to their iterative denoising process, but both classifier and classifier-free guidance strategies have been effectively adapted to offline MBO.
  • Flow matching shares many benefits with diffusion models but offers improved efficiency and performance.
  • EBMs are uniquely valuable in offline MBO for their ability to naturally penalize out-of-distribution samples and flexibly encode complex, domain-specific objectives. However, they are limited by slow MCMC-based sampling and unstable training dynamics.
  • GFlowNets, as a control framework, are particularly well-suited for offline MBO due to their ability to generate diverse, high-reward solutions and explore safely under uncertainty, though training cost is relatively high due to flow consistency constraints.
Practical Guide on Selecting Generative Models.
We have discussed various generative models. The choice of an appropriate model primarily depends on the data modality. Each type of data typically has a community-endorsed, state-of-the-art generative model. For instance, when the design is a prompt and the black-box property is the downstream task performance of an LLM given that prompt, autoregressive GPT-style models are the dominant choice due to strong empirical performances (Yang et al., 2024). Similarly, when the design is an image and the target property is a fine-grained visual attribute (e.g., a specific hairstyle), diffusion models are preferred for their superior image generation capabilities (Dhariwal & Nichol, 2021).
While model selection is primarily guided by data modality, this does not diminish the importance of discussing conditional generation strategies across different model families. In offline MBO, each generative model typically requires a tailored conditioning approach. For example, RGD (Chen et al., 2024) exploits the probability flow ODE of diffusion models to derive a refined proxy distribution and optimize the strength parameter—an approach unique to diffusion models. Similarly, MIN (Kumar & Levine, 2020) utilizes a GAN to directly map high-scoring property values to candidate designs through one-step generation, optimizing the values via a forward-inverse consistency loss, which is incompatible with the multi-step sampling process of diffusion models. Furthermore, certain models like VAEs support latent code manipulation, a feature generally unavailable in standard GANs due to the lack of an explicit latent representation.
There are cases where multiple generative models perform well for the same data modality. Protein sequence design is a representative example, and it is one of the most extensively studied tasks in offline MBO. In this setting, both autoregressive (AR) models (Elnaggar et al., 2021) and diffusion models (Wang et al., 2024) have demonstrated strong empirical performance. When deciding between the two, several additional considerations become important:
  • Likelihood estimation: Offline MBO often requires estimating the likelihood of candidate designs to detect and avoid out-of-distribution (OOD) samples. AR models can compute exact likelihoods by factorizing the joint distribution into conditional probabilities. In contrast, diffusion models rely on variational approximations such as the ELBO, which is computationally challenging. Thus, when precise likelihood computation and rigorous OOD detection are essential, AR models are generally preferred.
  • Sample diversity: Generating diverse samples is crucial for effective exploration in offline MBO. Diffusion models naturally provide broader mode coverage compared to AR models, making them better suited for tasks where diversity and exploration are prioritized.
  • Sample manipulation: The ability to perform localized edits on generated samples is critically important for control in offline MBO. Diffusion models support such fine-grained manipulation through their iterative denoising process. For instance, in antibody design, diffusion models can selectively generate CDR regions while preserving the antibody framework (Luo et al., 2022)—a capability not easily realized with AR models. Moreover, diffusion models benefit from their bidirectional nature (conditioning on both past and future context) and are compatible with advanced guided sampling techniques, enabling more flexible and nuanced control. A trade-off, however, is that diffusion models often have a fixed sequence length, whereas AR models natively support variable-length generation.
  • Availability of pretrained models: Practical deployment is often influenced by the accessibility of strong pretrained models. Fortunately, both AR (Elnaggar et al., 2021) and diffusion (Wang et al., 2024) models are publicly available and widely used. Leveraging these publicly available models can help mitigate OOD issues effectively.
  • Sampling efficiency: Efficient sampling is essential for scaling offline MBO to large candidate sets. AR models generate sequences token by token and are generally faster for shorter sequences. Diffusion models denoise entire sequences over multiple steps, which can be more efficient for longer sequences, though their overall speed depends on the number of denoising iterations used.
Summary: In offline MBO for protein sequence design, diffusion models may be more suitable due to their superior sample diversity and manipulability, despite limitations in exact likelihood estimation.